The ball of a float has a diameter of 4-in, and the thickness of the metal is 0.05 in. How much does it weigh if it is made of copper? (Copper weighs 556 lb. per cu. ft.)



1
Expert's answer
2021-06-11T06:10:48-0400

Volume of the entire ball"= \frac{4}{3} \pi ^3"

Volume "=\frac{4}3*\frac{22}7*2^3"

Volume "=33.524 in^3"

The volume of the hollow area of the ball "=\frac{4}3 \pi^3"

V"=\frac{4}3*\frac{22}7*[2-0.05]^3"

"V=31.0719 in^3"

Volume of copper "=33.524-31.072"

"V of \space copper =2.452in^3"

Density of Cu"=0.324 lb\/in^3"

Mass of Cu"=Density \space of \space Cu*Volume \space of \space Cu"

Mass of Cu"=0.324*2.452"

Mass of Cu"=0.794 lb"

Weight of copper "=0.794 *2.204*10"

Weight of copper"=17.510kg"